Divide and simplify
4sqrt a^3b^2/3sqrt ab
Please help, I would appreciate it

Answer by stanbon(75887)

Divide and simplify
4sqrt a^3b^2/3sqrt ab
--------------------------------
(4sqrt(a^3b^2))/(3sqrt(ab))
---
= (4ab*sqrt(a))/(3sqrt(ab))
---
= (4/3)(sqrt[a/ab])
---
= (4/3)sqrt(1/b)
Multiply numerator and denominator by sqrt(b) to get:
---
= (4/3)(sqrt(b))/b)
============================
Cheers,
Stan H.